HW 1 Ch 17 Winter 2019
Answer #1

is it 1.8 cm ?

Total Number of neurons passing through 1.8 cm is

N=5.6*1011*0.018 =1.008*1010 ions

Charge entering 1.6 cm length of axon is

Q=(1.6*10-19)(1.008*1010)

Q=1.6128*10-9C =1.6128 nC